OpenStudy (anonymous):

please help me multiply (w+5)(w-2)

OpenStudy (anonymous):

what grade

OpenStudy (trojanpoem):

Is w <- omega ?

OpenStudy (trojanpoem):

Complex numbers ?

OpenStudy (anonymous):

im a senior lol and all the question asks is too multiply (w+5)(w-2)

OpenStudy (trojanpoem):

Am I kidding ? if this were omega , then w + w^2 + 1 = 0

OpenStudy (trojanpoem):

Fine , if just multiplication multiply the W times ( w ) then (-2) then multiply +5 times (w) then ( -2 ) that's it

OpenStudy (anonymous):

so I would multiply (w) with (-2) and +5 times (w) and (-2) ???

OpenStudy (trojanpoem):

Yeah, so if you have ( w + y + u + i ) ( i + o + k ) you'd multiply w times i , o, k then y times i , o , k and so on.

OpenStudy (anonymous):

@TrojanPoem is correct I was gonna say the same thing lol xD

OpenStudy (anonymous):

i got 20 i don't think I'm right I'm usually really bad at math lol

OpenStudy (anonymous):

me too but ur correct tho hun

OpenStudy (trojanpoem):

(w+5)(w-2) w^2 - 2w +5w - 10 = 0 w^2 - 3 w - 10 = 0

OpenStudy (trojanpoem):

(w-5)(w+2) = 0 w = 5 w= -2

OpenStudy (trojanpoem):

if equal to 0

OpenStudy (anonymous):

I am…. well thats a first XD r u sure cuz i feel like I'm wrong

OpenStudy (anonymous):

yerp im sure sweet =)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

ur wlcm. always happy to help